# Schwinger dark matter production

### v1

@@ Abstract
Building on recently constructed inflationary vector dark matter production mechanisms as well as studies of magnetogenesis, we show that a dark Schwinger mechanism can generate the observed dark matter relic abundance. This mechanism can take place during inflation and/or after reheating and leads to a feebly interacting dark sector which is comprised of dark fermions and/or scalars which can be as light as $\sim 10\,MeV$ and as heavy as $\sim 10^{10}\,GeV$ depending on how the (unknown) Schwinger current redshifts after inflation. The background dark electric field is generated during inflation and can continue sourcing Schwinger pair production after reheating until the dark charged particles become non-relativistic. The dark matter can interact very weakly via the exchange of ultralight dark photons and has a power spectrum which is peaked at very small scales, thus evading potential isocurvature constraints. We emphasize that this mechanism is viable even near the conformal limit where (purely) gravitational particle production is negligible. Thus dark matter can be produced solely via the Schwinger effect even with a very weak electric force.
@@ I Introduction
The nature and production mechanism of dark matter is still unknown. Significant research effort has been put into candidates like weakly interacting massive particles, axions, sterile neutrinos, but we have seen no experimental evidence for them yet Bertone and Tait (2018). This motivates studying alternative candidates as well. One class which has received considerable attention in the recent past is that of dark photon dark matter, produced via non-thermal mechanisms Graham et al. (2016); Agrawal et al. (2020); Dror et al. (2019); Co et al. (2019); Bastero-Gil et al. (2019); Long and Wang (2019); McDermott and Witte (2020); Nakai et al. (2020); Ahmed et al. (2020); Kolb and Long (2021); Salehian et al. (2021); Bastero-Gil et al. (2022); Gorghetto et al. (2022); Sato et al. (2022); Redi and Tesi (2022); Barrie (2022) where the dark photon is associated with a dark $U(1)_{D}$ abelian gauge theory. At some point before matter radiation equality the dark photon must become massive, even if very light, in order to provide a cold (non-relativistic) dark matter candidate. However, it has been argued East and Huang (2022) that the formation of vortices in these scenarios spoils the possibility of the dark photon being dark matter [^1] However, these constraints can be softened in some dark photon dark matter scenarios Cyncynates and Weiner (2023). One can consider adding scalar and/or fermion fields charged under the $U(1)_{D}$ to the dark sector, of which the dark photon is the gauge boson. Utilizing one of the dark photon production mechanisms above, all that is needed is a way to transfer the energy density from the dark photon to the dark the ‘dark electrons’ [^2] We will generically refer to both scalars and fermions charged under the dark $U(1)_{D}$ as dark electrons. which can then constitute the dark matter.
@@
Such a ‘dark QED’ setup was studied in Arvanitaki et al. (2021) which utilized the mechanism of Graham et al. (2016) to generate a dark electric field composed of the longitudinal mode of the (massive) dark photon which is produced via inflationary fluctuations. While the modes are produced during inflation, the cold vector condensate, which can be treated as a classical dark electric field, only forms in the radiation dominated epoch when the vector momentum modes re-enter the horizon Graham et al. (2016). Furthermore, since the dark electric field which is produced is not very large compared to Hubble, this mechanism requires efficient transfer of the dark sector energy density from the dark photon to the dark electrons which requires the dark sector to thermalize in order to produce enough dark matter. It was shown Arvanitaki et al. (2021) how via various QED processes such as electromagnetic cascades, and plasma dynamics, as well as a subdominant contribution from Schwinger pair production, the dark sector can thermalize with itself while remaining decoupled from the visible thermal bath. In this case the amount of dark electrons today is determined by the interplay between the various thermalization processes in the dark sector and one obtains a relic abundance which matches the one observed for dark matter in the mass range from $50\,MeV$ to $30\,TeV$. In order to achieve thermalization the dark $U(1)_{D}$ gauge coupling cannot be too small, roughly $\gtrsim 10^{-3}$. Furthermore, as in Graham et al. (2016), the dark photon must be massive for the whole of its cosmological evolution including during inflation.
@@
We consider a similar dark QED scenario with the same field content as in Arvanitaki et al. (2021), but instead where the dark electric field is generated by the mechanism proposed in Bastero-Gil et al. (2019, 2022) which is based on an axion like coupling between the gauge field and the inflaton. Compared to the longitudinal mode production mechanism Graham et al. (2016) discussed above, the mechanism utilized here produces much larger coherent dark electric fields already during inflation via one of the transverse vector modes. Since it is the transverse mode which is produced, this mechanism also works in the massless dark photon case. Once formed these background fields can then source Schwinger pair production during and, in principle, after inflation.
@@
After reheating, the large electric fields continue to accelerate the dark electrons which have been pair produced. This imposes a stringent upper bound on the dark gauge coupling coming from the condition that the dark electrons become non-relativistic by the time of matter radiation equality. As we’ll see, this guarantees that we are always in the regime where the dark sector does not thermalize with itself. However, in this case the dark sector does not need to thermalize since, with the much larger electric fields, enough dark electrons can be produced solely via Schwinger pair production to reproduce the dark matter relic abundance. In the thermalization case discussed above, there is also Schwinger pair production taking place, but any initial seed of dark electrons coming from Scwhinger production is washed out by the more dominant thermalization effects Arvanitaki et al. (2021).
In the vector production mechanism considered in Graham et al. (2016); Arvanitaki et al. (2021) one can have large enough gauge couplings to thermalize, but yet not violate the non-relativistic condition. This is because in this case the dark electric field is much smaller so the electric force which accelerates the dark electrons is much weaker than in the mechanism considered here Bastero-Gil et al. (2019, 2022). This leads us to conclude that mechanisms which give weak dark electric fields can thermalize and in fact they must thermalize in order to produce enough dark electrons to constitute the dark matter. Mechanisms with stronger dark electric fields require much smaller gauge couplings to not violate the non-relativistic condition leading to no thermalization in the dark sector. In this case however the much larger dark electric field can generate the necessary amount of dark matter solely via the Schwinger effect.
@@
At the end of inflation and the onset of reheating, which we assume instantaneous, we have an energy density in the dark sector which is dominated by the dark electric field along with a tiny seed of Schwinger produced dark electrons. Schwinger production continues after reheating, but becomes inefficient once the dark electrons become non-relativistic. At that point the energy density in the Schwinger produced dark electrons is still smaller than that of the dark electric field. However, from then on the dark electrons begin redshifting like non-relativistic matter and come to dominate the dark sector energy density by the time of matter-radiation equality. Thus, the final relic abundance is set by the balance between Schwinger pair production and redshifting effects. Following this evolution we show that Schwinger produced dark electrons can reproduce the observed dark matter relic abundance for masses potentially as light as $\sim 10\,MeV$ and as heavy as $\sim 10^{10}\,GeV$ for Hubble scales at the end of inflation in the range $10^{8}-10^{14}\,GeV$. We expect the dark electrons to have a power spectrum peaked around the same scales as the dark photon. This was shown to be at very small scales Bastero-Gil et al. (2022), thus evading constraints on isocurvature coming from measurements of the CMB.
@@
Interestingly, and perhaps surprisingly, the dark Schwinger production mechanism is viable even in the weak dark electric field regime where backreaction effects can be neglected. This depends crucially on the fact that it occurs in an expanding background and is not possible in flat space where strong fields are required for efficient Schwinger production. This feature of the weak field regime in de-Sitter space was explored in various studies Fröb et al. (2014); Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018); Domcke and Mukaida (2018); Stahl (2019); Sobol et al. (2019); Domcke et al. (2020); Sobol et al. (2020) examining magnetogenesis. In these studies the conserved Schwinger current during inflation was computed for both scalars Banyeres et al. (2018) and fermions Hayashinaka et al. (2016). We adapt these results to study the evolution of the energy density of the produced dark sector particles averaged over space in order to estimate the parameter space where we can obtain the correct dark matter abundance.
@@
We also emphasize that conformal invariance must be broken during inflation to generate particles. In inflationary Schwinger production the conformal symmetry breaking comes from the background electric field (which can be generated in various ways). This breaking of conformal invariance ‘leaks’ into the matter sector via the Schwinger mechanism allowing for production of charged particles even in the absence of conformal breaking in the fermion and scalar sectors. In this way even massless fermions or conformally coupled massless scalars, which would not be produced by gravitational particle production due to a changing metric Ford (1987); Kuzmin and Tkachev (1999), can still be produced during inflation via the Schwinger effect.
Finally, we comment that part of our analysis relies on the Schwinger current after reheating, but to our knowledge there are no studies of Schwinger production in a Friedmann-Robertson-Walker (FRW) background which can be applied to epochs of radiation and matter domination. Thus, with the aim of simply gaining intuition and an estimate of the amount of dark particles which can be produced, in this study we assume that the electrical conductivity in the dark sector at the end of inflation either remains constant or redshifts at some rate after reheating. We leave a detailed study of the Schwinger mechanism in a generic FRW background and more precise estimate of the dark matter parameter space as well as power spectrum to ongoing work.
@@ II Dark Schwinger production
As we’ll see below, with the conductivity as input we can compute the energy density of the Schwinger produced charged particles during inflation in terms of the energy density of the dark $E$-field.
@@
Here $\langle E^{2}\rangle$ represents the spacial average of the magnitude squared of the dark electric field, We also have $n=3$ for the non-relativistic case and $n=4$ for the relativistic case while the overdot denotes a derivative with respect to cosmic time $t$. This evolution equation can can be derived from the energy momentum tensor and expresses energy conservation where $\langle E\cdot J\rangle$ acts as a source term for the $\chi$ energy density. A corresponding term, with the opposite sign, appears in the evolution equation for the energy density of the dark electric field. This must be sourced during inflation, as we discuss in the next section, but is depleted by the $-\sigma\langle E^{2}\rangle$ term, which is due to the Schwinger production  [^3] For these considerations to be sensible, $\langle E\cdot J\rangle$ must be positive, that is $\sigma>0$. It was pointed out in Hayashinaka et al. (2016) that the conductivity could turn negative in the limit of weak electric force and small mass of the charged particle. However this effect is unlikely to be physical Banyeres et al. (2018). We believe this point deserves further investigation which we leave for ongoing work. In this study we assume that the conductivity is always positive.. We work in the small conductivity regime where,
@@
where the energy density in the dark electric field is given by $\rho_{E}=\frac{1}{2}E^{2}$ Bastero-Gil et al. (2022). We see that at the end of inflation the energy density of the dark charged particles is proportional to the dark electric field energy density, but suppressed by the (dimensionless) conductivity. This result holds for both scalars and fermions and is independent of the source for the dark electric field. The only requirement is that it be (approximately) constant within a Hubble patch and contribute some amount to the energy density of the dark vector. Thus any mechanism which generates a constant background dark electric field during inflation will also Schwinger produce some amount of dark particles. If the dark sector does not thermalize with itself, the amount of dark charged particles today is determined simply by how much they are produced by the Schwinger effect and redshifting.
@@
In visible QED, after inflation and reheating the Standard Model (SM) plasma thermalizes and erases any initial seed of Schwinger produced electrons as well as any classical electric field generated during inflation. Similar conclusions follow for a dark QED sector which thermalizes with itself Arvanitaki et al. (2021). In this work we take the dark gauge coupling to be small enough that the dark QED sector never thermalizes within itself. We will see that for the dark electric field production mechanism considered here, this follows from requiring $\chi$ to be non-relativistic by matter radiation equality which imposes an upper bound on $g_{D}$ (see Eq. (49)). We also assume the dark sector is secluded from the visible one. Thus, any classical dark electric field will persist after reheating and redshift like radiation, all the while continuing to pair produce dark charged particles via the Schwinger mechanism.
After reheating, which we assume to be instantaneous for simplicity, we can write Eq. (5) as,
$\dot{\rho}_{\chi}+nH(t)\rho_{\chi}=\sigma(t)2\rho_{E}(t)\,,$ (10)
where we parametrize how various quantities redshift as,
$H(t)$ $=H_{\rm RH}\left(\frac{T}{T_{\rm RH}}\right)^{\delta}\,,$
$\sigma(t)$ $=\sigma_{\rm end}\left(\frac{a_{\rm end}}{a}\right)^{\alpha}=\sigma_{\rm RH}\left(\frac{T}{T_{\rm RH}}\right)^{\alpha}\,,$
$\rho_{E}(t)$ $=\rho_{E}^{\rm end}\left(\frac{a_{\rm end}}{a}\right)^{4}=\rho_{E}^{\rm RH}\left(\frac{T}{T_{\rm RH}}\right)^{4}\,,$ (11)
where RH refers to quantities defined at the end of reheating. In radiation domination one has $\delta=2$, but we keep it unspecified for now to allow for other possibilities ($\delta=3/2$ for matter domination for instance). It is unknown how the conductivity evolves after reheating as it requires calculation of the Schwinger current in an FRW background. Regardless, it is reasonable to assume it will not increase given that the electric field decreases due to redshifting. We thus expect $\sigma$ either to remain constant or to redshift at some rate. We parametrize this ignorance with the power scaling parameter $\alpha\geq 0$.
It is convenient to rewrite Eq. (10) switching variables from cosmic time $t$ to temperature $T$,
$\frac{d\rho_{\chi}}{dT}-\frac{n}{T}\rho_{\chi}=-2\bar{\sigma}_{\rm RH}\frac{\rho_{E}^{\rm RH}}{T_{\rm RH}}\left(\frac{T}{T_{\rm RH}}\right)^{3+\alpha-\delta}\,,$ (12)
where $\bar{\sigma}_{\rm RH}\equiv\sigma_{\rm RH}/H_{\rm RH}$. We take as the initial boundary condition at the end of reheating the amount of energy density produced at the end of inflation from Eq. (9),
$\rho_{\chi}(T_{\rm RH})\equiv\rho_{\chi}^{\rm RH}=\rho_{\chi}^{\rm end}=\frac{\bar{\sigma}_{\rm end}}{n}2\rho_{E}^{\rm end}=\frac{\bar{\sigma}_{\rm RH}}{n}2\rho_{E}^{\rm RH}\,.$ (13)
Then general the solution to Eq. (12) is given by,
[…]
-- $\frac{(4+\alpha-\delta)}{n}\left(\frac{T}{T_{\rm RH}}\right)^{n}\Big{]}\,.$
To gain some intuition, consider the case in which the produced $\chi$ particles are relativistic, the universe is radiation dominated (visible sector radiation), and the conductivity does not redshift which gives,
$n=4\,,\qquad\delta=2\,,\qquad\alpha=0\,.$ (15)
Then Eq. (II) becomes,
$\rho_{\chi}(T)$ $=\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left[\left(\frac{T}{T_{\rm RH}}\right)^{2}-\frac{1}{2}\left(\frac{T}{T_{\rm RH}}\right)^{4}\right]$
$\simeq\bar{\sigma}_{\rm RH}\rho_{E}^{\rm RH}\left(\frac{T}{T_{\rm RH}}\right)^{2}\,,$ (16)
where in the last equality we neglect the $T^{4}$ term compared to $T^{2}$ as it becomes less important as the universe expands and the temperature decreases. In the absence of Schwinger production $\chi$ would redshift like radiation as $\propto T^{4}$ after reheating. However, the persistence of the dark electric field and of Schwinger production contributes to $\rho_{\chi}$ thus making it redshift slower as $\propto T^{2}$.

### v2

@@ Abstract
Building on recently constructed inflationary vector dark matter production mechanisms as well as studies of magnetogenesis, we show that an inflationary dark Schwinger mechanism can generate the observed dark matter relic abundance for ‘dark electron’ masses as light as $\sim 0.1$ eV and as heavy as $10^{12}$ GeV. The dark matter can interact very weakly via the exchange of light dark photons with a power spectrum which is peaked at very small scales, thus evading isocurvature constraints. This mechanism is viable even when (purely) gravitational particle production is negligible. Thus dark matter can be produced solely via the Schwinger effect during inflation including for light masses.
@@ I Introduction
The nature and production mechanism of dark matter is still unknown. Significant research effort has been put into candidates like weakly interacting massive particles, axions, sterile neutrinos, but we have seen no experimental evidence for them yet Bertone and Tait (2018). This motivates studying alternative candidates as well. One class which has received considerable attention in the recent past is that of dark photon dark matter, produced via non-thermal mechanisms Graham et al. (2016); Agrawal et al. (2020); Dror et al. (2019); Co et al. (2019); Bastero-Gil et al. (2019); Long and Wang (2019); McDermott and Witte (2020); Nakai et al. (2020); Ahmed et al. (2020); Kolb and Long (2021); Salehian et al. (2021); Bastero-Gil et al. (2022); Gorghetto et al. (2022); Sato et al. (2022); Redi and Tesi (2022); Barrie (2022) where the dark photon is associated with a dark $U(1)_{D}$ abelian gauge theory. At some point before matter radiation equality the dark photon must become massive, even if very light, in order to provide a cold (non-relativistic) dark matter candidate. However, it has been argued East and Huang (2022) that the formation of vortices in these scenarios spoils the possibility of the dark photon being dark matter [^1] However, these constraints can be softened in some dark photon dark matter scenarios Cyncynates and Weiner (2023).. One can consider adding to the dark sector scalar and/or fermion fields charged under the $U(1)_{D}$, of which the dark photon is the gauge boson. Utilizing one of the dark photon production mechanisms above, all that is needed is a way to transfer the energy density from the dark photon to the ‘dark electrons’ [^2] We will generically refer to both scalars and fermions charged under the dark $U(1)_{D}$ as dark electrons. which can then constitute the dark matter.
@@
Such a ‘dark QED’ setup was studied in Arvanitaki et al. (2021) which utilized the mechanism of Graham et al. (2016) to generate a dark electric field composed of the longitudinal mode of the (massive) dark photon which is produced via inflationary fluctuations. While the modes are produced during inflation, the cold vector condensate, which can be treated as a classical dark electric field, only forms in the radiation dominated epoch when the vector momentum modes re-enter the horizon Graham et al. (2016). Since the dark electric field which is produced is not very large compared to the Hubble scale, this mechanism requires efficient transfer of the dark sector energy density from the dark photon to the dark electrons which requires the dark sector to thermalize in order to produce enough dark matter. It was shown Arvanitaki et al. (2021) how via various dark QED processes such as electromagnetic cascades, and plasma dynamics, as well as a subdominant contribution from Schwinger pair production, the dark sector can thermalize with itself while remaining decoupled from the visible thermal bath. In this case the amount of dark electrons today is determined by the interplay between the various thermalization processes in the dark sector and one obtains a relic abundance which matches the one observed for dark matter in the mass range from $50$ MeV to $30$ TeV. In order to achieve thermalization the dark $U(1)_{D}$ gauge coupling cannot be too small, roughly $\gtrsim 10^{-3}$. Furthermore, as in Graham et al. (2016), the dark photon must be massive for the whole of its cosmological evolution including during inflation.
@@
We consider a similar dark QED scenario with the same field content as in Arvanitaki et al. (2021), but instead where the dark electric field is generated by the mechanism proposed in Bastero-Gil et al. (2019, 2022) which is based on an axion like coupling between the gauge field and the inflaton. Compared to the longitudinal mode production mechanism Graham et al. (2016) discussed above, the mechanism utilized here produces much larger dark electric fields already during inflation via one of the transverse vector modes, thus it also works in the massless dark photon case. Once formed these background dark electric fields can then source Schwinger pair production of dark electrons during inflation.
@@
At the end of inflation and the onset of reheating, which we assume instantaneous, we have an energy density in the dark sector which is dominated by the dark electric field along with a tiny seed of Schwinger produced dark electrons. In the thermalization case discussed above, there is also Schwinger pair production, but any initial seed of dark electrons coming from Scwhinger production is washed out by the more dominant thermalization effects Arvanitaki et al. (2021). Here we consider the regime where the dark sector does not thermalize with itself which implies an upper bound on the dark gauge coupling.
In this case the only source of dark electrons is from inflationary dark Schwinger production. After reheating Schwinger production ceases and at some point the dark electrons begin redshifting like non-relativistic matter, while the electric field continues redshifting like radiation, eventually coming to dominate the dark sector energy density after matter-radiation equality. Thus, the relic abundance is set by the balance between inflationary Schwinger production and redshifting effects.
@@
Following this evolution we show that Schwinger produced dark electrons can reproduce the observed dark matter relic abundance for masses potentially as light as $\sim 0.1$ eV and as heavy as $\sim 10^{12}$ GeV. The dark gauge coupling can range from $\sim 10^{-10}-1$ while the Hubble scale at the end of inflation can be in the range $\sim(100-10^{13})$ GeV. We expect the dark electrons to have a power spectrum peaked around the same scales as the dark photon which is at very small scales Bastero-Gil et al. (2022), thus evading constraints on isocurvature coming from measurements of the CMB.
@@
The Schwinger effect in de-Sitter space has been explored in various studies Fröb et al. (2014); Kobayashi and Afshordi (2014); Hayashinaka et al. (2016); Banyeres et al. (2018); Domcke and Mukaida (2018); Stahl (2019); Sobol et al. (2019); Domcke et al. (2020); Sobol et al. (2020) in the context of magnetogenesis to explain the potential existence of primordial magnetic fields. In these studies the conserved Schwinger current during inflation was computed for both scalars Banyeres et al. (2018) and fermions Hayashinaka et al. (2016). We adapt these results to study the evolution of the energy density of the produced dark sector particles averaged over space in order to estimate the parameter space where we can obtain the correct dark matter abundance.
@@
We also emphasize that conformal invariance must be broken during inflation to generate particles. In inflationary Schwinger production the conformal symmetry breaking comes from the background electric field (which can be generated in various ways). This breaking of conformal invariance ‘leaks’ into the matter sector via the Schwinger mechanism allowing for production of charged particles even in the absence of conformal breaking in the fermion and scalar sectors. In this way even massless fermions or conformally coupled massless scalars, which would not be produced by purely gravitational particle production due to a changing metric Ford (1987); Kuzmin and Tkachev (1999), can still be produced during inflation via the Schwinger effect.
This paper is organized as follows. We first give a brief review of Schwinger production during inflation and compute the dark sector energy density at the end of inflation. We also review the electric field production mechanism based on an axion like coupling between the gauge field and the inflaton. We then track the cosmic evolution of the dark sector and compute the relic abundance today. Finally, we examine the viable dark matter parameter space of this mechanism before discussing potential future directions and concluding.
@@ II Dark Schwinger production
As we’ll see below, with the conductivity as input we can compute the energy density of the Schwinger produced charged particles during inflation in terms of the energy density of the dark electric field.
@@
Here $\langle E^{2}\rangle$ represents the spatial average of the magnitude squared of the dark electric field, We also have $n=3$ for the non-relativistic case and $n=4$ for the relativistic case while the overdot denotes a derivative with respect to cosmic time $t$. This evolution equation can be derived from the energy momentum tensor and expresses energy conservation where $\langle E\cdot J\rangle$ acts as a source term for the $\chi$ energy density. A corresponding term, with the opposite sign, appears in the evolution equation for the energy density of the dark electric field. This must be sourced during inflation, as we discuss in the next section, but is depleted by the $-\sigma\langle E^{2}\rangle$ term, which is due to the Schwinger production  [^3] For these considerations to be sensible, $\langle E\cdot J\rangle$ must be positive, that is $\sigma>0$. It was pointed out in Hayashinaka et al. (2016) that the conductivity could turn negative in the limit of weak electric force and small mass of the charged particle. However this effect is unlikely to be physical Banyeres et al. (2018). We believe this point deserves further investigation which we leave for ongoing work. In this study we assume that the conductivity is always positive.. We work in the small conductivity regime where,
@@
where the energy density in the dark electric field (derived from the energy momentum tensor) at the end of inflation is given by Bastero-Gil et al. (2022),
$\rho^{\rm{end}}_{E}=\frac{1}{2}E^{2}.$ (10)
@@
We see that at the end of inflation the energy density of the dark charged particles is proportional to the dark electric field energy density, but suppressed by the (dimensionless) conductivity. This result holds for both scalars and fermions and is independent of the source for the dark electric field. The only requirement is that it be (approximately) constant within a Hubble patch and contribute some amount to the energy density of the dark vector [^4] We will use ‘dark vector’ and ‘dark photon’ interchangeably.. Thus any mechanism which generates a constant background dark electric field during inflation will also Schwinger produce some amount of dark particles.
@@
In visible QED, after inflation and reheating the Standard Model (SM) plasma thermalizes and erases any initial seed of Schwinger produced electrons as well as any classical electric field generated during inflation. Similar conclusions follow for a dark QED sector which thermalizes with itself Arvanitaki et al. (2021). In this work we take the dark gauge coupling to be small enough that the dark QED sector never thermalizes within itself. We also assume the dark sector is secluded from the visible one. Thus the dark matter relic abundance is determined only by the inflationary Schwinger effect and redshifting.
Whether the electric field can source Schwinger production after inflation is unclear since a calculation of the conductivity during radiation domination in an FRW background does not exist. However, as we show in the Appendix, after reheating the electric field begins oscillating Bastero-Gil et al. (2022) and redshifting like radiation and no longer accelerates the dark electrons efficiently. We therefore expect Schwinger production to be suppressed after inflation though quantifying this requires further investigation. In this study we assume any Schwinger production after reheating is negligible.